Determine the polynomial function whose graph passes through the points (2, 4), (3,6) and (5,10). Also sketch the graph of the polynomial function. (Using Cramer’s Method).
1
Expert's answer
2020-12-21T18:07:09-0500

We have three points. So, we shall use second degree polynomial fitting.


"p(x)=ax^2+bx+c"

For "(2,4)" ,

"4=a(2^2)+b(2)+c\\
4=4a+2b+c......................(1)"

For "(3,6),"

"6=a(3^2)+b(3)+c\\
6=9a+3b+c......................(2)"

For "(5,10),"

"10=a(5^2)+b(5)+c\\
10=25a+5b+c...................(3)"

Bring the three equations together.

"4a+2b+c=4\\
9a+3b+c=6\\
25a+5b+c=10"

"\begin{pmatrix}
 4 & 2 &1\\
 9 &3 &1\\
25 &5 &1\\
\end{pmatrix}\begin{pmatrix}a\\b\\c\\\end{pmatrix}=\begin{pmatrix}4\\6\\10\\\end{pmatrix}"

Using Cramer's rule.


"\Delta=\begin{vmatrix}
 4 & 2 &1\\
 9 &3 &1\\
25 &5 &1\\
\end{vmatrix}=-6"


"\Delta_x=\begin{vmatrix}
 4 & 2 &1\\
 6 &3 &1\\
10 &5 &1\\
\end{vmatrix}=0\\
a=\frac{\Delta_x}{\Delta}=\frac{0}{-6}=0"


"\Delta_y=\begin{vmatrix}
 4 & 4 &1\\
 9 &6 &1\\
25 &10 &1\\
\end{vmatrix}=-12\\
b=\frac{\Delta_y}{\Delta}=\frac{-12}{-6}=2"


"\Delta_z=\begin{vmatrix}
 4 & 2 &4\\
 9 &3 &6\\
25 &5 &10\\
\end{vmatrix}=0\\
c=\frac{\Delta_z}{\Delta}=\frac{0}{-6}=0"


So,

"p(x)=0.x^2+2.x+0\\
p(x)=2x"


Here is the graph of "p(x)=2x"